OpenStudy (anonymous):

Given log7=0.8451 and log2=0.3010 find the value of log28

OpenStudy (anonymous):

\(28=2^2\times 7\) is a start

OpenStudy (unklerhaukus):

\[\log(28)=\log(4\times7)=\]

OpenStudy (unklerhaukus):

\[\log ab=\log a+\log b\] \[\log c^n=n\log c\]

OpenStudy (anonymous):

log(28)=log(28)?

OpenStudy (unklerhaukus):

that is step zero @Sukhvir , can you see the next steps as well...

OpenStudy (anonymous):

log((28)(28))

OpenStudy (anonymous):

log((784))

OpenStudy (anonymous):

log(784) = 2.8943

myininaya (myininaya):

You need to use the rules mentioned by @UnkleRhaukus here @Sukhvir Factor 28 Then use the first property of logs mentioned by @UnkleRhaukus Also @satellite73 has already factored 28 for you \[28=2^2*7\] So \[\log(28)=\log(2^2 * 7)\] Now apply that first property here: \[\log(2^2 * 7)=\]

OpenStudy (anonymous):

what is the first property? :S

OpenStudy (unklerhaukus):

that the logarithm of a product of factors is the sum of the logarithm of the factors \[\log (a\times b)=\log (a)+\log (b)\]

OpenStudy (anonymous):

loh(8)+log(7)

OpenStudy (anonymous):

3log(2)+log(7)

OpenStudy (unklerhaukus):

\[\text{lol}(7\times8)=\text{lol}({56)}\]

OpenStudy (unklerhaukus):

\[\log(28)=\]

OpenStudy (anonymous):

ugh i cant figure this out can you show me? my math skills are not strong

OpenStudy (unklerhaukus):

do you know how to factorise 28 ?

OpenStudy (anonymous):

4*7

OpenStudy (anonymous):

log(28)=log(4)+log(7)

OpenStudy (unklerhaukus):

good, can you also factorise 4

OpenStudy (anonymous):

2*2

OpenStudy (unklerhaukus):

which can also be written (in index form) as....

OpenStudy (anonymous):

2log(2)?

OpenStudy (unklerhaukus):

a*a = \(a\times a\) = a^2 = \(a^2\)

OpenStudy (unklerhaukus):

4 = 2*2 = 2 ^ ?

OpenStudy (anonymous):

2^2

OpenStudy (unklerhaukus):

right , and the second rule of logs i wrote was the log of a number raised to an index is the index multiplied by the log of the number \[\log c^n=n\log c\] so \[\log(28)=\log(4)+\log(7)\]\[\quad\qquad=\log(2^2)+\log(7)\]\[\qquad\quad=\]\[\qquad\quad=\]

OpenStudy (anonymous):

log(28)=1.748

OpenStudy (unklerhaukus):

nope

OpenStudy (anonymous):

where did i go wrong?

OpenStudy (anonymous):

i did log of 2^3 and then log7 and added them

OpenStudy (unklerhaukus):

if you dont want to learn the second rule thats fine just apply the first rule again \[\log(28)=\log(4)+\log(7)=\log(2)+\log(2)+\log(7)\]

OpenStudy (unklerhaukus):

where did you get a three?

OpenStudy (anonymous):

i missed read the power

OpenStudy (unklerhaukus):

ah well \[\log 2^3+\log7=1.748\] so the rest of your method worked

OpenStudy (anonymous):

aha so i got 1.447 = log(28)?

OpenStudy (unklerhaukus):

thats good \[\large\color{red}\checkmark\]